# Advantage Alignment Algorithms

### v1

@@ A.3 Gradient of LOQA
$=\hat{\pi}^{2}(b_{t}|s_{t})\left(\nabla_{\theta_{1}}Q^{2}_{t}(b_{t})-\sum_{b}\hat{\pi}^{2}(b|s_{t})\nabla_{\theta_{1}}Q^{2}_{t}(b)\right)$ (33)
Where we used the quotient rule in line (30) and equation (4) in line(33). By the chain rule, the gradient of the log probability is:
@@
A.4 Proof of Theorem 2
@@
Differentiating (34) leads to a simplified gradient:
@@
$=\nabla_{\theta_{1}}\exp Q^{2}_{t}(b_{t})\frac{\sum_{b\neq b_{t}}\exp Q^{2}_{t}(b)}{\left(\exp Q^{2}_{t}(b_{t})+\sum_{b\neq b_{t}}\exp Q^{2}_{t}(b)\right)^{2}}$ (37)
$=\exp Q^{2}_{t}(b_{t})\nabla_{\theta_{1}}Q^{2}_{t}(b_{t})\frac{\sum_{b\neq b_{t}}\exp Q^{2}_{t}(b)}{\left(\exp Q^{2}_{t}(b_{t})+\sum_{b\neq b_{t}}\exp Q^{2}_{t}(b)\right)^{2}}$ (38)
@@
The difference between LOQA and Advantage Alignment lies in the extra scaling factor $(1-\tilde{\pi}^{2}(b_{t}|s_{t}))$ which accounts for the partition function. Plugging (41) into the generalized policy gradient equation (6) proves the theorem. ∎
A.5 Advantage Alignment Implementation
To more clearly see the Advantage Alignment formula as an influence over each indicual log probability term recall the formulation:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{k}A^{1}(s_{t},a_{t},b_{t})A^{2}(s_{k},a_{k},b_{k})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\right]$ (8)
Reordering the summations we get:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\sum_{k<t}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right]$ (42)
@@ A.5 Advantage Alignment Implementation
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\frac{1}{1+t}\left(\sum_{k<t}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right]$ (43)
Which accounts to increasing the probability of the actions that align the sum of the past advantages of the agent up to the current time-step $t-1$ and the advantage of the opponent at the current time-step, $t$. In our implementation we use equation (43), as it considers more terms in the future and works better in practice.
A.6 Proximal Advantage Alignment
@@
     Compute agent actor loss, $L_{a}^{1}$, summing equations (9) and (45)
     Compute opponent actor loss, $L_{a}^{2}$, summing equations (9) and (45)
@@
  end for
@@
We then apply it to the advantage alignment formula (8):
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{k}\min\left\{r_{n}(\theta_{1})A^{1}_{t}A^{2}_{k},\;\text{clip}\left(r_{n}(\theta_{1});1-\epsilon,1+\epsilon\right)A^{1}_{t}A^{2}_{k}\right\}\right]$ (9)
Where we denote $\pi^{1}_{n}(a_{t}|s_{t})$ to be the updated policy and $r_{n}(\theta_{1})=\pi^{1}_{n}(a_{t}|s_{t})/\pi^{1}(a_{t}|s_{t})$ is the ratio between the updated policy and the old policy. We used generalized advantage estimation (GAE) (Schulman et al.,, 2018) to compute the advantages in this expression.
@@ B.1 Iterated Prisoner’s Dilemma
We use an MLP layer connected to a GRU followed by another MLP head for both the actor and critic networks, similar to the architecture used in POLA (Zhao et al.,, 2022). We also use a replay buffer of agents collected during training, following Aghajohari et al., 2024b . All of our experiments run in 50 minutes in a nvidia A100 gpu.
@@ B.2 Coin Game
We use the same architecture used for IPD with an MLP connected to a GRU unit, followed by another MLP. We experimented with both AdAlign (Equation (8)) and Proximal AdAlign (Equation (9)), with AdAlign performing better (this is the one we used). All of our experiments run in 30 minutes in a nvidia A100 gpu.
@@ B.3 Negotiation Game
Hyperparameters: Please refer to 4 for the hyperparameters used in our negotiation game experiments. We use a replay buffer on our gather trajectories although the rate that it is mixed with fresh trajectories is small.
@@
Note that the optimization of the agents in the negotiation game is unstable, preventing us from taking the last checkpoint. In our experiments in Fig 4(a) we select the checkpoint that corresponds to the best achieved return for the agent during the optimization of the agent and the opponent. While we are not completely certain, we observe the instability happens when the policy distribution concentrates around the maximum possible proposal which is $5$. We clipped the $\operatorname{atanh}$ operation in our implementation for more numerical stability. All of our experiments run in 1 hour on a nvidia A100 gpu.
@@ C.1 Negotiation Game Training Curves
Figure 5 shows the training curves of advantage alignments on 10 seeds.
@@ C.2 Coin Game Full League Results
Figure 6: The head-to-head results of all variants of the cion game agents experimented with in this paper. All numbers are an average of 10 seeds of one type of agent with 10 seeds of another type of agent, where each pair play 32 games.

### v3

@@ A.3 Gradient of LOQA
$=\hat{\pi}^{2}(b_{t}|s_{t})\left(\nabla_{\theta_{1}}Q^{2}_{t}(b_{t})-\sum_{b}\hat{\pi}^{2}(b|s_{t})\nabla_{\theta_{1}}Q^{2}_{t}(b)\right),$ (45)
where we used the quotient rule in line (42) and equation equation 4 in line(45). By the chain rule, the gradient of the log probability is:
@@
A.4 Gradient of LOQA in Continuous Action Spaces
We derive the gradient of the opponent’s policy $\pi^{2}(b|s)$ with respect to the agent’s parameters $\theta_{1}$, assuming a continuous action space.
The opponent’s policy is defined as:
[…]
where $Q^{2}(s,b)$ is the Q-value of the opponent for action $b$, and $\mathcal{A}$ is the continuous action space.
Our goal is to compute the gradient of $\pi^{2}(b|s)$ with respect to $\theta_{1}$, the parameters of agent 1, which affect $Q^{2}(s,b)$ through interactions.
Taking the log of $\pi^{2}(b|s)$, we get:
[…]
The gradient of $\log\pi^{2}(b|s)$ with respect to $\theta_{1}$ is:
[…]
which simplifies to:
[…]
Now, applying the chain rule to compute the gradient of $\pi^{2}(b|s)$, we get:
[…]
The Integral above is intractable, which makes continuous action LOQA hard to scale.
A.5 Proof of Theorem 2
Proof.
@@ A.5 Proof of Theorem 2
Differentiating equation 53 leads to a simplified gradient:
@@
The difference between LOQA and Advantage Alignment lies in the extra scaling factor $(1-\tilde{\pi}^{2}(b_{t}|s_{t}))$ which accounts for the partition function. Plugging equation 58 into the generalized policy gradient equation equation 6 proves the theorem. ∎
A.6 Advantage Alignment Implementation
To more clearly see the Advantage Alignment formula as an influence over each individual log probability term recall the formulation:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t+1}\left(\sum_{k<t}\gamma^{t-k}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right].$ (8)
@@ A.6 Advantage Alignment Implementation
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\frac{1}{1+t}\left(\sum_{k<t}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right].$ (59)
Which accounts to increasing the probability of the actions that align the sum of the past advantages of the agent up to the current time-step $t-1$ and the advantage of the opponent at the current time-step, $t$. In our implementation we use equation 59, as it considers more terms in the future and works better in practice.
A.7 Proximal Advantage Alignment
@@
     Compute agent actor loss, $L_{a}^{1}$, using equation 9
     Compute opponent actor loss, $L_{a}^{2}$, using equation 9
@@
We can combine the two policy gradient terms into a single one to come up with a proximal Advantage Alignment formulation:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}_{t}\nabla_{\theta^{1}}\log\pi^{1}(a_{t}|s_{t})+\beta\gamma\sum_{t=0}^{\infty}\gamma^{t}\left(\sum_{k<t}\gamma^{t-k}A^{1}_{k}\right)A^{2}_{t}\nabla_{\theta^{1}}\log\pi^{1}(a_{t}|s_{t})\right]$ (60)
Where $\beta$ is the weight put into the Advantage Alignment loss (the negative inverse of the Boltzmann constant times the temperature). Then we have:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(A^{1}_{t}+\beta\gamma\left(\sum_{k<t}\gamma^{t-k}A^{1}_{k}\right)A^{2}_{t}\right)\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right].$ (61)
This is just the normal advantage policy gradient with a modified advantage $A^{*}$:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{*}_{t}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right],\text{ where }A^{*}_{t}=A^{1}_{t}+\beta\gamma\left(\sum_{k<t}\gamma^{t-k}A^{1}_{k}\right)A^{2}_{t}.$ (62)
@@
Now we apply it to the Advantage Alignment formulation that uses the modified advantage on the policy gradient equation 62:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\min\left\{r_{n}(\theta_{1})A^{*}_{t},\;\text{clip}\left(r_{n}(\theta_{1});1-\epsilon,1+\epsilon\right)A^{*}_{t}\right\}\right],$ (9)
where we denote $\pi^{1}_{n}(a_{t}|s_{t})$ to be the updated policy and $r_{n}(\theta_{1})=\pi^{1}_{n}(a_{t}|s_{t})/\pi^{1}(a_{t}|s_{t})$ is the ratio between the updated policy and the old policy. We used generalized advantage estimation (GAE) (Schulman et al.,, 2018) to compute the advantages in this expression. Algorithm 2 summarizes the implementation of Proximal Advantage Alignment.
A.8 Proof of Theorem 3
Let $\theta_{1},...,\theta_{n}$ be the parameter each agent, $\pi_{\theta_{i}}(a|s)$ be the policies represented by those parameters, and $V_{i}(\theta_{1},...,\theta_{n})$ be the value function of agent $i$ as a function of all the other agents.
Lemma 2 (Zero Advantages At Nash).
For all Nash Equilibria of the game, if there exist parameters $\theta_{1}^{*},...,\theta_{n}^{*}$ such that $\pi_{\theta_{i}^{*}}=\pi_{i}^{*}$, where $\pi_{i}^{*}$ is the policy of agent $i$ at the Nash, then for all action-state pairs with non-zero probability under the Nash policies, we have $A_{i}(a|s)=0$.
Proof.
By the Bellman Optimality Equation, at an optimal policy the value of agent $i$ becomes $V_{i}^{*}(s)=\arg\max_{a}Q^{*}(a,s)$, hence all actions with non-zero probability under $\pi_{i}^{*}$ have the same $Q^{*}(a,s)$, and since $A(a,s)\equiv Q(a,s)-V(s)$, the advantage will vanish. ∎
We now use lemma 2 to prove that the Advantage Alignment term is zero at a Nash equilibrium.
Proof.
Under Advantage Alignment, the updates we take can be represented by
$\theta^{\prime}_{i}$ $\leftarrow\theta_{i}+\alpha\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{i},\pi^{-i}}}\left[\sum_{t=0}^{\infty}B^{i}_{t}\>\nabla_{\theta_{i}}\log\pi_{\theta_{i}}(a_{t}|s_{t})\right]$ (65)
$B^{i}_{t}$ $\equiv A^{i}_{t}+\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{i},\pi^{-i}}}\left[\sum_{j\neq i}\left(\sum_{k<t}\gamma^{t-k}A^{i}_{k}\right)A^{j}_{t}\right]$ (66)
But by lemma 2, $A^{j}(b_{t}|s_{t})=0$ for all actions at a Nash, hence the second term vanishes, as does the first term for the same reason. ∎
A.9 N-Player Advantage Alignment
Consider the $n$-player setup, we can use the policy gradient theorem and assumption 2 to derive the following expression:
$\nabla_{\theta_{1}}V^{i}(\mu)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{i},\pi^{-i}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{i}_{t}\left(\nabla_{\theta_{i}}\log\pi^{i}(a^{i}_{t}|s_{t})+\sum_{j\neq i}\nabla_{\theta_{i}}\log\pi^{j}(a^{j}_{t}|s_{t})\right)\right]$ (67)
Which naturally leads to the following modification to the PPO advantage following the derivation used in Proximal Advantage Alignment:
${A^{i}_{t}}^{*}=A^{i}_{t}+\beta\gamma\left(\sum_{k<t}\gamma^{t-k}A^{i}_{k}\right)\sum_{j\neq i}A^{j}_{t}$ (68)
Here we use the standard game theory notation of $i$ to refer to the current player and $-i$ to refer to all other players. Similarly $a^{i}_{t}$ denotes the action of player $i$ at time $t$.
@@ B.1 Iterated Prisoner’s Dilemma
We use an MLP layer connected to a GRU followed by another MLP head for both the actor and critic networks, similar to the architecture used in POLA (Zhao et al.,, 2022). We also use a replay buffer of agents collected during training, following Aghajohari et al., 2024b . All of our IPD experiments run in 50 minutes in a nvidia A100 gpu.
@@ B.2 Coin Game
We use the same architecture used for IPD with an MLP connected to a GRU unit, followed by another MLP. We experimented with both Advantage Alignment (Equation equation 8) and Proximal Advantage Alignment (Equation equation 9), with Advantage Alignment performing better (this is the one we used). All of our Coin Game experiments run in 30 minutes in a nvidia A100 gpu.
@@ B.3 Negotiation Game
We experimented with both Advantage Alignment (Equation equation 8) and Proximal Advantage Alignment (Equation equation 9), with the original Advantage Alignment performing better.
@@
Hyperparameters: Please refer to 3 for the hyperparameters used in our negotiation game experiments. We use a replay buffer on our gather trajectories although the rate that it is mixed with fresh trajectories is small.
@@
Note that the optimization of the agents in the negotiation game is unstable, preventing us from taking the last checkpoint. In our experiments in Fig 3(a) we select the checkpoint that corresponds to the best achieved return for the agent during the optimization of the agent and the opponent. While we are not completely certain, we observe the instability happens when the policy distribution concentrates around the maximum possible proposal which is $5$. We clipped the $\operatorname{atanh}$ operation in our implementation for more numerical stability. All of our Negotiation Game experiments run in 1 hour on a nvidia A100 gpu.
B.4 Melting Pot’s Commons Harvest Open
We experimented with both Advantage Alignment (Equation equation 8) and Proximal Advantage Alignment (Equation equation 9), with Proximal Advantage Alignment performing better.
Agent’s Architecture: In the Commons Harvest Open environment, agents receive observations consisting of a local view of the environment in the form of raw pixel data. Each observation is an image frame capturing the agent’s immediate surroundings. We use a 3 layer convolutional neural network, following (Mnih et al.,, 2013), to encode the observations, which are then passed to a GTrXL tranformer (Parisotto et al.,, 2019).
Encoder: The observation frames are processed by an encoder. The encoder is a GTrXL transformer network (Parisotto et al.,, 2019). The GTrXL network consists of 3 transformer layers, each with a model dimension of 192 and a feedforward dimension of 192. The transformer is capable of handling sequences up to a maximum length of 1000 steps, capturing temporal dependencies in the agents’ observations. In practice, we use a context length of 15.
Critic: The output of the encoder is then fed to a two-layer Multi-Layer Perceptron (MLP) with ReLU non-linearities for the critic module of the agent. To provide temporal context, we concatenate the current time step to the encoder’s output before feeding it to the critic. This helps the critic estimate the value of the state more accurately, as the remaining time in an episode can affect the expected return.
Actor: The actor network shares the encoder with the critic. The output of the encoder is passed through another MLP with ReLU non-linearities to produce logits over discrete action choices. The policy is modeled as a categorical distribution over these actions, which include turning around, moving in different directions, and zapping other agents.
Hyperparameters: Please refer to Table 4 for the hyperparameters used in our Commons Harvest Open experiments.
Table 4: Commons Harvest Open Hyperparameters
Parameter Value
Self-Play True
Batch Size 512
Optimizer (Actor) Learning Rate $1\times 10^{-5}$
Optimizer (Critic) Learning Rate $1\times 10^{-5}$
Entropy Beta 0.1
Advantage Alignment Weight 1.0
Clip Gradient Norm 10.0
Transformer Layers 3
Transformer Model Dimension 192
Transformer Feedforward Dimension 192
Discount Factor ($\gamma$) 0.99
PPO Clip Range 0.1
PPO Updates per Batch 2
Normalize Advantages True
Context Length 15
We use a parallelized environment with 6 copies of Commons Harvest Open to make training more efficient. Following LOQA (Aghajohari et al., 2024b, ), we keep a replay buffer of past agent parameters to ensure robustness against a distribution of policies. From this replay buffer we sample 2 agents at each iteration and play against 5 self-play agents with the current version of the policy. For each environment, we use the 5 on-policy trajectories to compute losses for the actor and critic. In total, our Commons Harvest Open experiments last 24 hours on an nvidia L40s gpu.
@@ C.1 Negotiation Game Training Curves
Figure 6 shows the training curves of Advantage Alignment on 10 seeds.
@@ C.2 Coin Game Full League Results
Figure 7: The head-to-head results of all variants of the coin game agents experimented with in this paper. All numbers are an average of 10 seeds of one type of agent with 10 seeds of another type of agent, where each pair play 32 games. We ablate Advantage Alignment masking different components of the gradient. Cooperative (C), masks when both advantages are positive; Empathetic (E), masks when the advantage of the agent is positive and the advantage of the opponent is negative; Vengeful (V), masks when the advantage of the agent is negative and the advantage of the opponent is positive; Spiteful (S), masks when both advantages are negative.
C.3 Ablation Study Commons Harvest Open
Figure 8: Sample trajectories for Proximal Advantage Alignment agents with different $\beta$ weight. We select the best of 10 seeds for each value of $\beta$. On the first row: $\beta=0.5$, agents reach a policy where they try to consume the apples as fast as possible. On the second row: $\beta=1$, agents reach a ”bush guarding” policy, zapping any other agents coming into the same bush. On the third row: $\beta=2$, agents reach a policy where they rotate around specific paths, preventing the extinction of the bushes.
Interestingly the value of $\beta$, which is used to control the weight of the advantage alignment term in equation 10, leads agents to converge to different policies. With a low value of the weight ($\beta=0.5$), we empirically observed that most runs converge to a greedy policy that attempts to consume apples as soon as possible. With a value of $\beta=1$, we find policies that show a ”bush guarding” behavior preventing other agents from approaching their bush, and consuming apples within that bush with moderate restraint. This is the policy that shows the best evaluation performance in figure 4. With high values of the weight ($\beta=2$), most runs find a rotating strategy in which agents stick to eating only a subset of the apples on each bush. This policy has the highest pro-social return out of all of them. However, the rotating strategy is also vulnerable to exploitation from greedy agents and does poorly in the evaluation scenarios. Figure 8 showcases what these policies look like in practice.